which property of real numbers are shown
-6(4x+5)=-24x-30

Answer by stanbon(75887)

which property of real numbers are shown
-6(4x+5)=-24x-30
-----
Distribution Law::
Multiplication by -6 is distributed of addition of 4x and 5
-6(4x+5 = -6*(4x) + -6(5) = -24x - 30
----------------
Cheers,
Stan H.
-------------